OpenStudy (anonymous):

Solve the equation. |x + 5| = 9 A. {–14, 4} B. {–14, –4} C. {4, 14} D. {–4, 14}

OpenStudy (ranga):

|A| = B means A = +B or A = -B |x + 5| = 9 means x + 5 = 9 or x + 5 = -9 x = 9 - 5 = 4 x = -9 - 5 = -14 x = { -14, 4 }